# Learning Color Equivariant Representations

### v4

@@ 1 Introduction
The tremendous progress of image classification in the last decade can be readily attributed to the development of deep convolutional neural networks (Krizhevsky et al., 2012; Simonyan & Zisserman, 2014; He et al., 2016). The highly nonlinear mapping and large parameter space does not lend itself to interpretation easily, but even in early networks, representations of geometry and color were observed and recognized for their importance (Krizhevsky et al., 2012; Lenc & Vedaldi, 2015). While a relatively large body of literature has worked to improve the robustness of convolutional neural networks to geometric transformations (Bruna & Mallat, 2013; Cohen & Welling, 2016; Gilmer et al., 2017; Qi et al., 2017; Hinton et al., 2018; Greydanus et al., 2019), improving their robustness to perceptual variation has garnered considerably less attention.
@@
Figure 1: Color-equivariant network. (a) The equivariance of our hue-equivariant model is illustrated by the commutativity of the (hue) rotation and neural network mapping. A hue rotation of 90 in the input image space (top-left to bottom-left), results in a feature map rotation at each layer of the network (top-right to bottom-right). Corresponding feature maps are highlighted with a blue border. (b) An input image (left) is lifted to the hue-saturation group (right) by shifting its hue and saturation values. For comparison, we illustrate the CEConv lifting layer in Appendix B.1.
@@ 2 Related Work
Group convolutional networks have also been designed for use in the context of scale symmetry. In Esteves et al. (2018), the authors represent input images in log-polar coordinates where 2D rotation and scale transformations present as 2D translations. To navigate the semi-group structure of scale transformations, the authors in Worrall & Welling (2019) approximate the scale space as finite and use dilated convolutions in a group equivariant architecture. The SREN network proposed in Sun & Blu (2022) designs for equivariance to the continuous 2D rotation and isotropic scaling group by constraining filter representations to be the linear combination of a windowed Fourier basis.
@@ 4 Method
Hue group and group action. In the HSL color space, hue is represented by angular position, and can therefore be identified with the 2D rotation group. As in group convolutional network (Cohen & Welling, 2016), we consider a finite group representation. Specifically, we identify elements of the discretized hue group, $H_{N}$, where the subscript $N$ indicates the order of (i.e. the number of elements in) the group, with those of the cyclic group $C_{N}$.
@@
The action of $\varphi_{h}$ on an input image, and $\phi_{h}$ on a feature map are shown in Figure 6(a). We prove $\varphi_{h}$ and $\phi_{h}$ are group actions in Appendix A.
Saturation group and group action. We expand the notion of color equivariance proposed in Lengyel et al. (2024) to include equivariance to saturation shifts. In the HSL color space, saturation can be represented by a real number in the interval $[0,1]$; we introduce two approximations to give the saturation space group structure. First, we observe that there is a bijection between the real numbers and the open interval $(0,1)$, so we can use the structure of the group $(\mathbb{R},+)$. Second, we consider a finite subset of the group, a necessary and commonly used practice in both conventional CNNs (recall that the translation group is infinite) and group-equivariant CNNs such as Worrall & Welling (2019). With these approximations, we identify elements of the discretized saturation group $S_{N}$, with those of the integers with addition, $(\mathbb{Z},+)$.
@@
The action of $\varphi_{s}$ on an input image, and $\phi_{s}$ on a feature map are shown in Figure 2. We prove $\varphi_{s}$ and $\phi_{s}$ are group actions in Appendix A. We define the luminance group and group action similarly (see Appendix E.1).
@@
We prove $\varphi_{hs}$ and $\phi_{hs}$ are group actions in Appendix A.
@@
Here we denote an element of the product space $g_{x,i}$, where the subscript $x$ references an element of $\mathbb{Z}^{2}$, and $j$ references the element $h_{j}$ in $H_{N}$. The input image $f^{0}$ and lifting filters $\psi_{m,n}$ are functions on $\mathbb{Z}^{2}$. This is the approach taken in CEonv (Lengyel et al., 2024). The authors shift the hue of a filter in the RGB space by rotating its values about an axis passing through the point $p=(1,1,1)$ as shown in Figure 8(b) in Appendix B.1. As described in Lengyel et al. (2024), this approach results in invalid hue rotations for all discretizations of the hue group that are not symmetries of the axis-aligned RGB cube (i.e., $N=1$ and $N=3$). To remedy this, the authors project invalid RGB values to the nearest point on the RGB cube. However, this projection breaks invertibility of the hue shift (see Figure 3), yielding a descriptor that is only approximately hue equivariant.
@@
Equivariance of the lifting layer. We can show that our lifting layer is equivariant to hue shifts. We let $f^{1}(g_{x,j})=\varphi_{h}(h_{j},f^{0})(x)$ as in Equation 10 and show that a hue shift of $f^{0}$ results in a hue shift of $f^{1}$. In the first step we use the fact that $\varphi_{h}$ is a group action (see Appendix A.1), and in the second step we use commutativity of the hue group:
@@
where $\psi^{l}$ is a function on $G$. Group convolution on the saturation and hue-saturation groups are performed analogously.
@@ 5 Experiments
In this section, we highlight the sample efficiency of our model in the context of hue generalization; the stability of our hue-equivariant representations compared to CEConv (Lengyel et al., 2024); the utility of a notion of color equivariance that includes saturation and luminance; and the utility of our representations for color-based sorting. Our models achieve strong performance on extensive experiments, against competitive baselines. Additional experiments are presented in Appendix E.
@@ 5.1 Hue-shift MNIST Classification
We demonstrate improved generalization to global hue-shifts and higher sample efficiency compared to a conventional CNN model with a similar number of parameters on our Hue-shift MNIST dataset. Our dataset is a variation of MNIST (LeCun et al., 1998) with 60k training examples and 10k test examples classified into one of 10 categories. Additional details are provided in Appendix C.1.
@@ 5.2 Hue-shift 3D Shapes Classification
We demonstrate improved generalization to local hue-shifts and significantly reduced equivariance error compared to CEConv on the 3D Shapes classification dataset (Burgess & Kim, 2018). The 3D Shapes dataset consists of RGB images of 3D shapes, where the color of the shape, the floor, and the walls vary across examples, as does the scale and orientation of the shape. There are 48k examples in the training set, and 12k examples in each of the test sets. Additional details are provided in Appendix C.2.
@@ 5.3 Camelyon17 Classification: Saturation shift in the wild
We demonstrate improved generalization to saturation-shifts compared to ResNet50 (He et al., 2016) and CEConv-3 on the Camelyon17 classification dataset (Bandi et al., 2018). The Camelyon17 dataset consists of images of human tissue collected from five different hospitals. Variation in tissue images results from variation in the data collection and processing procedures. The dataset consists of 387,490 examples, 302,436 of which are used for training and 85,054 of which are used for testing. Hospitals 1, 2, and 3 are represented in the training set, and hospital 5 is represented in the testing set. Additional details are provided in Appendix D.3.
@@ 5.4 Color shift in the wild
We demonstrate improved generalization to global color-shifts compared to ResNet (He et al., 2016) and CEConv on the Caltech-101 (Li et al., 2022), Oxford-IIT Pets (Parkhi et al., ), Stanford Cars (Krause et al., 2013), STL-10 (Coates et al., 2011), CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009) datasets. We report results on the ImageNet dataset in Appendix E.2 and provide additional details for all experiments in Appendix D.
@@
Figure 7: Color sorting on CIFAR-10. We show images from CIFAR-10 automobile class ordered by pairwise distance using Hue-4 and ResNet44 feature maps. The structure of the Hue-4 feature maps naturally allow for color based sorting, whereas the ResNet44 feature maps do not. We include visualizations of the entire automobile class sorted using Hue-$N$ and CEConv-$N$ in Appendix B.2.
@@ Acknowledgments
We sincerely thank the ICLR 2025 reviewers and area chair for their thoughtful feedback; the discussions and additional experiments have meaningfully improved our paper.
@@ A Hue and saturation group action
In this section we prove the proposed hue group action and saturation group action satisfy the axioms for a group action. By definition of a group action, $\varphi_{h}$ is a group action if it satisfies the following properties:

### v6

@@ 1 Introduction
The tremendous progress of image classification in the last decade can be readily attributed to the development of deep convolutional neural networks (Krizhevsky et al., 2012; Simonyan & Zisserman, 2014; He et al., 2016). The highly nonlinear mapping and large parameter space does not lend itself to interpretation easily, but even in early networks, representations of geometry and color were observed and recognized for their importance (Krizhevsky et al., 2012; Lenc & Vedaldi, 2015). While a large body of literature has worked to improve the robustness of neural networks to geometric transformations (Bruna & Mallat, 2013; Cohen & Welling, 2016; Gilmer et al., 2017; Qi et al., 2017; Hinton et al., 2018; Esteves et al., 2018a; Greydanus et al., 2019; Zhong & Allen-Blanchette, 2025), improving their robustness to perceptual variation has garnered considerably less attention.
@@
Figure 1: Color-equivariant network. (a) The equivariance of our hue-equivariant model is illustrated by the commutativity of the (hue) rotation and neural network mapping. A hue rotation of 90 in the input image space (top-left to bottom-left), results in a feature map rotation at each layer of the network (top-right to bottom-right). Corresponding feature maps are highlighted with a blue border. (b) An input image (left) is lifted to the hue-saturation group (right) by shifting its hue and saturation values. For comparison, we illustrate the CEConv lifting layer in Appendix C.1.
@@ 2 Related Work
Group convolutional networks have also been designed for use in the context of scale symmetry. In Esteves et al. (2018b), the authors represent input images in log-polar coordinates where 2D rotation and scale transformations present as 2D translations. To navigate the semi-group structure of scale transformations, the authors in Worrall & Welling (2019) approximate the scale space as finite and use dilated convolutions in a group equivariant architecture. The SREN network proposed in Sun & Blu (2022) designs for equivariance to the continuous 2D rotation and isotropic scaling group by constraining filter representations to be the linear combination of a windowed Fourier basis.
@@ 4 Method
Hue group and group action. In the HSL color space, hue is represented by angular position, and can therefore be identified with the 2D rotation group. As in group convolutional network (Cohen & Welling, 2016), we consider a finite group representation. Specifically, we identify elements of the discretized hue group, $H_{N}$, where the subscript $N$ indicates the order of (i.e. the number of elements in) the group, with those of the cyclic group $C_{N}$. We prove $H_{N}$ is a group in Appendix A.1.
@@
The action of $\varphi_{h}$ on an input image, and $\phi_{h}$ on a feature map are shown in Figure 6(a). We prove $\varphi_{h}$ and $\phi_{h}$ are group actions in Appendix B.1.
Saturation group and group action. We expand the notion of color equivariance proposed in Lengyel et al. (2024) to include equivariance to saturation shifts. In the HSL color space, saturation can be represented by a real number in the interval $[0,1]$; we introduce two approximations to give the saturation space group structure. First, we observe that there is a bijection between the real numbers and the open interval $(0,1)$, so we can use the structure of the group $(\mathbb{R},+)$. Second, we consider a finite subset of the group, a necessary and commonly used practice in both conventional CNNs (recall that the translation group is infinite) and group-equivariant CNNs such as Worrall & Welling (2019). With these approximations, the discretized saturation group $S_{N}$ is isomorphic to the integers with addition, $(\mathbb{Z},+)$. We prove $S_{N}$ is a group in Appendix A.2.
@@
The action of $\varphi_{s}$ on an input image, and $\phi_{s}$ on a feature map are shown in Figure 2. We prove $\varphi_{s}$ and $\phi_{s}$ are group actions in Appendix B.2. We define the luminance group and group action similarly (see Appendix F.1).
@@
We prove $\varphi_{hs}$ and $\phi_{hs}$ are group actions in Appendix B.
@@
Here we denote an element of the product space $g_{x,i}$, where the subscript $x$ references an element of $\mathbb{Z}^{2}$, and $j$ references the element $h_{j}$ in $H_{N}$. The input image $f^{0}$ and lifting filters $\psi_{m,n}$ are functions on $\mathbb{Z}^{2}$. This is the approach taken in CEonv (Lengyel et al., 2024). The authors shift the hue of a filter in the RGB space by rotating its values about an axis passing through the point $p=(1,1,1)$ as shown in Figure 8(b) in Appendix C.1. As described in Lengyel et al. (2024), this approach results in invalid hue rotations for all discretizations of the hue group that are not symmetries of the axis-aligned RGB cube (i.e., $N=1$ and $N=3$). To remedy this, the authors project invalid RGB values to the nearest point on the RGB cube. However, this projection breaks invertibility of the hue shift (see Figure 3), yielding a descriptor that is only approximately hue equivariant.
@@
Equivariance of the lifting layer. We can show that our lifting layer is equivariant to hue shifts. We let $f^{1}(g_{x,j})=\varphi_{h}(h_{j},f^{0})(x)$ as in Equation (10) and show that a hue shift of $f^{0}$ results in a hue shift of $f^{1}$. In the first step we use the fact that $\varphi_{h}$ is a group action (see Appendix B.1), and in the second step we use commutativity of the hue group:
@@
where $\psi^{l}$ is a function on $G$. Group convolution on the saturation, luminance, and hue-saturation-luminance groups are performed analogously.
@@ 5 Experiments
In this section, we highlight the sample efficiency of our model in the context of hue generalization; the stability of our hue-equivariant representations compared to CEConv (Lengyel et al., 2024); the utility of a notion of color equivariance that includes saturation and luminance; and the utility of our representations for color-based sorting. Our models achieve strong performance on extensive experiments, against competitive baselines. Additional experiments are presented in Appendix F.
@@ 5.1 Hue-shift MNIST Classification
We demonstrate improved generalization to global hue-shifts and higher sample efficiency compared to a conventional CNN model with a similar number of parameters on our Hue-shift MNIST dataset. Our dataset is a variation of MNIST (LeCun et al., 1998) with 60k training examples and 10k test examples classified into one of 10 categories. Additional details are provided in Appendix D.1.
@@ 5.2 Hue-shift 3D Shapes Classification
We demonstrate improved generalization to local hue-shifts and significantly reduced equivariance error compared to CEConv on the 3D Shapes classification dataset (Burgess & Kim, 2018). The 3D Shapes dataset consists of RGB images of 3D shapes, where the color of the shape, the floor, and the walls vary across examples, as does the scale and orientation of the shape. There are 48k examples in the training set, and 12k examples in each of the test sets. Additional details are provided in Appendix D.2.
@@ 5.3 Camelyon17 Classification: Saturation shift in the wild
We demonstrate improved generalization to saturation-shifts compared to ResNet50 (He et al., 2016) and CEConv-3 on the Camelyon17 classification dataset (Bandi et al., 2018). The Camelyon17 dataset consists of images of human tissue collected from five different hospitals. Variation in tissue images results from variation in the data collection and processing procedures. The dataset consists of 387,490 examples, 302,436 of which are used for training and 85,054 of which are used for testing. Hospitals 1, 2, and 3 are represented in the training set, and hospital 5 is represented in the testing set. Additional details are provided in Appendix E.3.
@@ 5.4 Color shift in the wild
We demonstrate improved generalization to global color-shifts compared to ResNet (He et al., 2016) and CEConv on the Caltech-101 (Li et al., 2022), Oxford-IIT Pets (Parkhi et al., ), Stanford Cars (Krause et al., 2013), STL-10 (Coates et al., 2011), CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009) datasets. We report results on the ImageNet dataset in Appendix F.2 and provide additional details for all experiments in Appendix E.
@@
Figure 7: Color sorting on CIFAR-10. We show images from CIFAR-10 automobile class ordered by pairwise distance using Hue-4 and ResNet44 feature maps. The structure of the Hue-4 feature maps naturally allow for color based sorting, whereas the ResNet44 feature maps do not. We include visualizations of the entire automobile class sorted using Hue-$N$ and CEConv-$N$ in Appendix C.3.
@@ Acknowledgments
We sincerely thank the ICLR 2025 reviewers and area chair for their thoughtful feedback; the discussions and additional experiments have meaningfully improved our paper. We thank Minseok Kim for their insights and discussion.
@@ A Hue and Saturation Group
In this section we prove the proposed discretized hue group and saturation group satisfy the axioms for a group. By definition, a group is a non-empty set $G$ along with a binary operation on $G$ that satisfies:
  1. 1.
    Closure. For all $a,b\in G$,
    $a\cdot b\in G.$ (18)
  2. 2.
    Associativity. For all $a,b,c\in G$,
    $\left(a\cdot b\right)\cdot c=a\cdot\left(b\cdot c\right).$ (19)
  3. 3.
    Identity element. There exists an element $1\in G$ so that for all $a\in G$,
    $a\cdot 1=1\cdot a=a.$ (20)
  4. 4.
    Inverse element. For each $a\in G$, there exists $a^{-1}\in G$ so that,
    $a\cdot a^{-1}=a^{-1}\cdot a=1.$ (21)
A.1 Hue Group
We show the set $H_{N}$,
$H_{N}=\left\{0,\ldots,\nicefrac{{2\pi k}}{{N}},\ldots,\nicefrac{{2\pi(N-1)}}{{N}}\right\},$ (22)
where $0\leq k<N$, together with the binary operation, $\cdot:H_{N}\times H_{N}\to H_{N}$, where
[…]
for any $a,b\in H_{N}$, is a group. To do this, we prove the group axioms (18)-(21).
Closure is satisfied if for any $a,b\in H_{N}$, we have $a\cdot b\in H_{N}$. Any $a,b\in H_{N}$ can be expressed $a=\nicefrac{{2\pi k_{a}}}{{N}}$ and $b=\nicefrac{{2\pi k_{b}}}{{N}}$ for some integers $k_{a}$ and $k_{b}$ in the range $0\leq k_{a},k_{b}<N$. The product of $a$ and $b$ can then be written
[…]
To prove closure, we consider two cases: $k_{a}+k_{b}<N$, and $k_{a}+k_{b}\geq N$. In the first case, that is, $k_{a}+k_{b}<N$, we immediately have
$a\cdot b=\frac{2\pi\left(k_{a}+k_{b}\right)}{N}\in H_{N},$ (25)
since $k_{a}+k_{b}$ is in the range $0\leq k_{a}+k_{b}<N$. In the second case, that is, $k_{a}+k_{b}\geq N$, we can write $k_{a}+k_{b}$ in quotient remainder form,
$k_{m}*N+k_{r}=k_{a}+k_{b},$ (26)
where $k_{m}$ is the divisor, and $k_{r}$ is the remainder. Using this form we can write the product of elements $a$ and $b$
[…]
$=\frac{2\pi k_{r}}{N}.$ (29)
Since $0\leq k_{r}<N$, we have that $a\cdot b\in H_{N}$, and closure is satisfied.
Associativity is satisfied if for any $a,b,c\in H_{N}$, we have $a\cdot(b\cdot c)=(a\cdot b)\cdot c$. For any $a,b,c\in H_{N}$, we can write $a=\nicefrac{{2\pi k_{a}}}{{N}}$, $b=\nicefrac{{2\pi k_{b}}}{{N}}$, and $c=\nicefrac{{2\pi k_{c}}}{{N}}$ for some integers $k_{a}$, $k_{b}$ and $k_{c}$ in the range $0\leq k_{a},k_{b},k_{c}<N$. Then we have
[…]
Since modulo distributes over addition, the above can be written
[…]
Using the identity property of modulo (i.e., $(a\;\mathrm{mod}\;b)(\mathrm{mod}\,b)=a\;\mathrm{mod}\;b$), we can write
[…]
Using the distributive property of modulo, we can write
[…]
$=a\cdot(b\cdot c),$ (36)
and associativity is satisfied.
Identity element. An element $e\in H_{N}$ is an identity element if for every $a\in H_{N}$, we have $e\cdot a=a\cdot e=a$. Consider the element $0\in H_{N}$, then for every $a=\nicefrac{{2\pi k_{a}}}{{N}}\in H_{N}$ we have
[…]
By commutativity of addition, we have $0\cdot a=a$, and 00 is the identity element of $H_{N}$.
Inverse element. There exists an inverse for every element of the group if for every $a\in H_{N}$ there exists an element $a^{-1}\in H_{N}$ so that $a\cdot a^{-1}=a^{-1}\cdot a=e$, where $e\in H_{N}$ is the identity element. We can write any element $a\in H_{N}$, in the form $a=\nicefrac{{2\pi k_{a}}}{{N}}$ where $k_{a}$ is an integer in the range $0\leq k_{a}<N$. An inverse element $a^{-1}$ of $a$ satisfying $a\cdot a^{-1}=a^{-1}\cdot a=0$ can be written in the form $a^{-1}=\nicefrac{{2\pi\left(N-k_{a}\right)}}{{N}}$,
[…]
By commutativity of addition we have $a^{-1}\cdot a=0$. It remains to show that $a^{-1}$ is in the set $H_{N}$. Since $k_{a}$ is an integer in the range $0\leq k_{a}<N$, then $N-k_{a}$ is in the range $0<N-k_{a}\leq N$, and therefore, $a^{-1}=\nicefrac{{2\pi\left(N-k_{a}\right)}}{{N}}$ is in the set $H_{N}$ and existence of an inverse is satisfied.
[…]
A.2 Saturation Group
We show that the set $S_{N}$
$S_{N}=\left\{\ldots,\nicefrac{{-2k}}{{N}},\nicefrac{{-k}}{{N}},\nicefrac{{0}}{{N}},\nicefrac{{k}}{{N}},\nicefrac{{2k}}{{N}},\ldots\right\},$ (39)
generated by $\nicefrac{{k}}{{N}}$ with $k$ integer valued, together with the binary operation $\cdot:S_{N}\times S_{N}\to S_{N}$, where
$a\cdot b\mapsto a+b$ (40)
for any $a,b\in S_{N}$ is a group. To do this, we prove the group axioms (18)-(21).
Closure is satisfied if for any $a,b\in S_{N}$, we have $a\cdot b\in S_{N}$. Any $a,b\in S_{N}$ can be expressed $a=\nicefrac{{m_{a}*k}}{{N}}$ and $b=\nicefrac{{m_{b}*k}}{{N}}$ for some integers $m_{a}$ and $m_{b}$. The product of $a$ and $b$ can then be written
$a\cdot b=\frac{m_{a}*k}{N}+\frac{m_{b}*k}{N}=\frac{(m_{a}+m_{b})*k}{N}.$ (41)
Since $m=m_{a}+m_{b}$ is an integer (the set of integers is closed under addition), we have that $a\cdot b=\nicefrac{{m*k}}{{N}}$, is in the set $S_{N}$, and, closure is satisfied.
Associativity is satisfied if for any $a,b,c\in S_{N}$, we have $a\cdot(b\cdot c)=(a\cdot b)\cdot c$. For any $a,b,c\in S_{N}$, we can write $a=\nicefrac{{m_{a}*k}}{{N}}$, $b=\nicefrac{{m_{b}*k}}{{N}}$, and $c=\nicefrac{{m_{c}*k}}{{N}}$ for some integers $m_{a}$, $m_{b}$, and $m_{c}$. Then we can write,
$\left(a\cdot b\right)\cdot c=\left(\frac{m_{a}*k}{N}+\frac{m_{b}*k}{N}\right)+\frac{m_{c}*k}{N}=\frac{(m_{a}+m_{b}+m_{c})*k}{N}$ (42)
and similarly,
$c\cdot\left(b\cdot c\right)=\frac{m_{a}*k}{N}+\left(\frac{m_{b}*k}{N}+\frac{m_{c}*k}{N}\right)=\frac{(m_{a}+m_{b}+m_{c})*k}{N}.$ (43)
Therefore, $a\cdot(b\cdot c)=(a\cdot b)\cdot c$, and associativity is satisfied.
Identity element. An element $e\in S_{N}$ is an identity element if for every $a\in S_{N}$, we have $e\cdot a=a\cdot e=a$. Consider $0\in S_{N}$, then for every $a\in S_{N}$ where $a=\nicefrac{{m_{a}*k}}{{N}}\in S_{N}$ we have,
$a\cdot 0=\frac{m_{a}*k}{N}+\frac{0*k}{N}=a.$ (44)
By commutativity of addition, we have $0\cdot a=a$, and 00 is the identity element of $S_{N}$.
Inverse element. There exists an inverse for every element of the group if for every $a\in S_{N}$ there exists an element $a^{-1}\in S_{N}$ so that $a\cdot a^{-1}=a^{-1}\cdot a=e$, where $e\in S_{N}$ is the identity element. We can write any element $a\in S_{N}$, in the form $a=\nicefrac{{m_{a}*k}}{{N}}$ where $m_{a}$ is integer valued. An inverse element $a^{-1}$ of $a$ satisfying $a\cdot a^{-1}=a^{-1}\cdot a=0$ can be written in the form $a^{-1}=\nicefrac{{-m_{a}*k}}{{N}}$,
$a\cdot a^{-1}=\frac{m_{a}*k}{N}+\frac{-m_{a}*k}{N}=0.$ (45)
By commutativity of addition we have $a^{-1}\cdot a=0$. It remains to show that $a^{-1}$ is in the set $S_{N}$. Since $-m_{a}$ is an integer, $a^{-1}=\nicefrac{{-m_{a}*k}}{{N}}$ is in the set $S_{N}$ and existence of an inverse is satisfied.
Having shown that the set $S_{N}$ together with the binary operation $(a\cdot b)\mapsto a+b$ satisfies all four axioms of a group, we have that $S_{N}$ is in fact a group.
B Hue and saturation group action
In this section we prove the proposed hue group action and saturation group action satisfy the axioms for a group action. By definition of a group action, $\varphi$ is a group action if it satisfies the following properties: